Find Z if the area to the right of Z is 0.67

A)     0.9100

B)     0.0421

C)     -0.4500

D)     -0.390
Answer #1

SOLUTION:

Using standard normal table,

P(Z > z) =0.67

= 1 - P(Z < z) = 0.67  

= P(Z < z) = 1 - 0.67

= P(Z < z ) = 0.33

z = -0.4500 ( using z table )